OpenStudy (anonymous):

Find all real solutions to equation:x^4/3-11x^2/3+24=0

OpenStudy (anonymous):

solve \[u^2-11u+24=0\] then replace u by \[x^{\frac{2}{3}}\]

OpenStudy (sasogeek):

\[\large (\frac{x^4}{3-11}) (\frac{x^2}{3+24})=0 \text{ is this the question?} \]

OpenStudy (anonymous):

once I factor, I get (u-8)(u-3)?

OpenStudy (anonymous):

\[x^{\frac43}-11x^{\frac23}+24=0\]

OpenStudy (sasogeek):

\[\large x^{4/3}-11x^{2/3}+24=0??? \]

OpenStudy (sasogeek):

ohhhh

OpenStudy (anonymous):

\[x^{4/3}-11x ^{2/3}+24=0\]

OpenStudy (sasogeek):

then what satellite said :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what does \[x ^{2/3}=8\] equal for x?

OpenStudy (anonymous):

no, 3/2

OpenStudy (sasogeek):

that means \[\large \sqrt[3]{x^2}=8 \]

OpenStudy (sasogeek):

you raise both sides to the power 3, that will clear the cube root on the x and make 8^3=512 then you find the square root of 512, and that is your x :)